# High-Sensitivity Photonic Crystal Biosensors using Topological Light Trapping

### v1

@@ I Introduction
In our PBG-based biosensors, the sensitivity is defined as the rate of change of the resonant frequency with the increment of the analyte thickness for fixed analyte refractive index or the increment of analyte refractive index for fixed analyte thickness. The limit-of-detection is defined as the minimum change in analyte coating thickness (or refractive index) required to optically detect that a change occurred. For detection based on an optical transmission resonance, the limit-of-detection is usually determined by the quality factor of the resonance. The frequency shift of the resonance must be at least the full-width at half-maximum of the original resonance for unambiguous detection. Much effort has been made to achieve a good biosensor with both a high sensitivity and a low limit-of-detectionFarida25 . Photonic crystal biosensors provide a unique opportunity to simultaneously achieve high sensitivity and low-limit-of-detection using a minimal volume of bio-fluid.
@@ II.2 Quality factor enhancement
The topological light-trapping mode for the structure depicted in Fig. 2(c) has a quality factor of only 1520.1.
@@
The quality factor can be improved by two independent methods. The first involves moving the resonance towards the middle of the PBG. This decreases the spatial extent of the localized optical mode and its coupling to glass encasing structure. By reducing the size of the rectangular dielectric block in the defect region from $(0.2a\times 0.4a)$ to $(0.12a\times 0.24a)$, the optical resonance moves to $a/\lambda=0.279794$, almost in the middle of the band gap. The quality factor is improved from $1520.1$ to $2299.7$. The second method to decrease the coupling of the localized mode to the glass encasing is by widening the entire flow channel. By additionally increasing the number of unit cells on each side of the defect from $l/a=5.5$ to $l/a=6.5$, the quality factor is further improved to $9021.9$, as shown in Fig. 4. The quality factor, in fact, rises exponentially with increasing $l/a$ until it saturates due to weak absorption of light in a real medium. PC-book08 Enlarging the chip width increases the numerical simulation time greatly. Therefore, we stop at a reasonable flow channel half-width ($l/a=5.5$ for most of our simulations).
@@ II.3 Sensitivity enhancement
Analyte-binding on surfaces of only one square block within the defect region, as in Ref. 13, provides a resonant frequency shift of only $0.0006$ $2\pi c/a]$ per analyte-thickness $t=0.1a$.
@@ III Two topological defects: mode hybridization and Spectral independence
For biosensors capable of multi-parametric detection, which can distinguish different concentrations of multiple disease markers in a single measurement, multiple defects are required. This gives rise to mode hybridization, which complicates the interpretion of transmission spectra due to correlated spectral responses from different analyte-bindings. We first study a chip with two defects to analyze mode hybridization and its relation to transmission-levels. We analyze correlations in frequency shifts of the two separate resonances, and how to achieve spectral independence.
@@ III.3 Intermediate Hybridization
We functionalize this chip for biosensing and numerically simulate its transmission spectra for the various analyte-binding cases, as shown in Fig. 10(b)-(d). Due to increased mode hybridization, the spectral responses of the two resonances become more correlated. Nevertheless, different binding cases are clearly distinguishable expecially if we monitor the transmission-level changes, as shown in the Truth Table (Fig. 10(e)). The frequency-shifts by themselves are also distinct. For example, with L-binding, the left-like resonance $|\text{L}\rangle^{\prime}$ will red-shift, inducing a correlated red-shift of the right-like resonance $|\text{R}\rangle^{\prime}$. Since the field strength of the left-like mode is higher than that of the right-like mode in the left-defect region, the left-like resonance $|\text{L}\rangle^{\prime}$ will red-shift more. The resonant frequency difference $\Delta\omega$ becomes larger with L-binding. This reduces hybridization and makes the transmission-levels of both resonances decline. Similar analysis applies to the other binding cases.
@@ V Conclusion and Outlook
Even higher sensitivity can be achieved with more elaborate designs or other types of defects. For example, increasing the number of rectangular strips within one unit cell from eight to twelve, with properly chosen strip width, such as $w_{def}=0.02a$, will improve the sensitivity further to 0.00758 $2\pi c/a]$ frequency shift per analyte-thickness $t=0.05a$, which is $46\%$ higher than our optimal sensitivity here. However, such a chip may be more challenging to fabricate. The chips given here all involve topological defects that involve a disruption of lattice periodicity. These were combined with substitutional defects (involving a shape change within a unit cell) to enhance sensitivity and adjust mode frequencies. Purely substitutional defect with fragmented rectangular strips can also achieve a high sensitivity. However, such chips typically require period doubling along the direction of fluid flow to enable optical coupling by an external plane wave.Abdullah15 ; Shuai16 ; Abdullah19 ; Dragan21 Combining topological domain-walls with substitutional defects enables additional design flexibility, not possible with purely substitutional defects. This may be useful in the design of 3D chips with finite pillar height. Nevertheless, other issues will arise. For example, an improperly fragmented unit cell may radiate more light into the 3rd dimension, thereby lowering the quality factor. Accordingly, further design considerations may be required to retain high sensitivity while not not losing light vertically. We hope that our design guidelines described herein will aid the development of real-world 3D photonic crystal biosensors, requiring minimal biofluid volume, while achieving low limit-of-detection, high sensitivity, and spectral independence, with efficacious transmission quality.

### v2

@@ I Introduction
In our PBG-based biosensors, the sensitivity is defined as the rate of change of the resonant frequency with the increment of the analyte thickness for fixed analyte refractive index or the increment of analyte refractive index for fixed analyte thickness. The limit-of-detection is defined as the minimum change in analyte coating thickness (or refractive index) required to optically detect that a change occurred. For detection based on an optical transmission resonance, the limit-of-detection is usually determined by the quality factor (Q) of the resonance. The frequency shift of the resonance must be at least the full-width at half-maximum of the original resonance for unambiguous detection. Much effort has been made to achieve a good biosensor with both a high sensitivity and a low limit-of-detectionFarida25 . Photonic crystal biosensors provide a unique opportunity to simultaneously achieve high sensitivity and low-limit-of-detection using a minimal volume of bio-fluid.
Other mechanisms for biosensing have also been proposed in the literature. For example, nanoplasmonic devices have demonstrated single-molecule sensitivity through extreme electromagnetic field enhancements in nanoscale hot spotsTaylor2017 ; Wang2013 ; Mauriz2021 ; Radziuk2015 . However, the purpose of point-of-care diagnostics is to provide a nearly instantaneous (in a few minutes) outcome, without the need for more time-consuming testing in an external laboratory. The long timescale required for a single molecule to diffuse within a bio-fluid sample and find a specific nanometer scale hotspot makes the nanoplasmonic device less suitable for point-of-care medical diagnostics. In other words, it may require many hours (or longer) for the nanoscale “mouse” in a realistic fluid sample to find the nanoscale “mousetrap”. This might be overcome by having a high concentration of the disease-marker molecules in the sample, but then it would not be truly “single molecule detection” (since a large number of molecules was necessary for timely detection). Alternatively, the timescale problem could be overcome by a device hosting a large number of spatially separated nanoplasmonic hotspots, creating a large spectral density of nanoplasmonic resonances. Within this “noisy” background it may be difficult to identify the spectral shift of a single resonance due to the attachment of a single molecule to the single hotspot. Plasmonic resonances are intrinsically broad and lossy (low Q), and have limited reproducibility, significant photothermal heating, and restricted multiplexing capabilityVollmer2008 ; Jin2025 .
Our lab-in-a-PC enables nearly instantaneous detection and diagnosis by providing a spatially broad “net”, occupying a sizeable volume of photonic crystal flow channel, for the capture of disease-marker moleculesAbdullah15 ; Shuai16 ; Abdullah19 ; Dragan21 . Each “line-defect” within the 2D PC extends along the length of the fluid flow direction. The resulting optical modes, offer much higher quality factors and narrower line-widths than nanoplasmonic devices, leading to high refractive-index figures of merit and quantitative, label-free detection. These high-Q photonic modes enable precise spectral readout and low limits-of-detection, while maintaining stability and low absorption losses. Our larger mode volumes (enabling rapid detection) and weaker local field enhancements make unassisted single-molecule detection more challenging than for plasmonic hot-spots. Nevertheless, our photonic platform provides spectral precision, reproducibility, and rapid detection.
@@ II Light trapping by a domain wall
Note: Here we use a typical value $n=1.35$ ($\varepsilon=1.8225$) for the refractive-index, corresponding to human blood serum. The choice of specimen may shift the exact numerical values of the resonant frequencies, but the overall detection scheme will remain the same. For example, a fluid specimen with a higher refractive index will cause all the resonant frequencies to shift slightly lower. It is a simple matter to re-calibrate the device to account for such a change in specimen. For all analytes (disease biomarkers) that bind to the functionalized silicon surfaces of the optical biosensor, a representative refractive index of $n=1.45$ ($\varepsilon=2.1025$) is adopted. This analyte index applies to all figures where spectral shift as a function of analyte layer thickness is depicted. If an alternative analyte with a higher or lower refractive index (from a different specimen) is chosen, the corresponding spectral shifts will accordingly be higher or lower.
@@ II.2 Quality factor enhancement
The topological light-trapping mode for the structure depicted in Fig. 2(c) has a quality factor of only 1520.1. The full-width at half-maximum (FWHM) of this resonance is $\Delta\omega^{\text{FWHM}}=\omega/Q=1.7\times 10^{-4}[2\pi c/a]$.
@@
The quality factor can be improved by two independent methods. The first involves moving the resonance towards the middle of the PBG. This decreases the spatial extent of the localized optical mode and its coupling to glass encasing structure. By reducing the size of the rectangular dielectric block in the defect region from $(0.2a\times 0.4a)$ to $(0.12a\times 0.24a)$, the optical resonance moves to $a/\lambda=0.279794$, almost in the middle of the band gap. The quality factor is improved from $1520.1$ to $2299.7$. The second method to decrease the coupling of the localized mode to the glass encasing is by widening the entire flow channel. By additionally increasing the number of unit cells on each side of the defect from $l/a=5.5$ to $l/a=6.5$, the quality factor is further improved to $9021.9$, as shown in Fig. 4. The FWHM, which is inversely proportional to the quality factor, is reduced to $\Delta\omega^{\text{FWHM}}=\omega/Q=3.1\times 10^{-5}[2\pi c/a]$.
The quality factor, in fact, rises exponentially with increasing $l/a$ until it saturates due to weak absorption of light in a real medium. PC-book08 The structures we illustrate have widths (in the light propagation direction) ranging from 11 to 13 lattice constants. This corresponds to about 3-4 optical wavelengths. On such short length scales, absorption is generally negligible. A very small residual absorption would place a limit on the highest quality factor we can obtain for any of our optical resonances. In the absence of absorption, the quality factors increase exponentially with the width of photonic crystal chip (because the optical mode amplitude decreases exponentially with distance from the center of each defect). If the residual absorption is characterized by an effective absorption length scale, the quality factor of the optical transmission resonance will reach its maximum value (saturate) once the chip width equals the absorption length. That is to say, there is no advantage in making a photonic crystal chip that is wider than the absorption length scale.
From a computational perspective, enlarging the chip width increases the numerical simulation time greatly. Therefore, we stop at a reasonable flow channel half-width ($l/a=5.5$ for most of our simulations).
@@ II.3 Sensitivity enhancement
Analyte-binding on surfaces of only one square block within the defect region, as in Ref. 13, provides a resonant frequency shift of only $0.0006$ $2\pi c/a]$ per analyte-thickness $t=0.1a$. With such a low sensitivity, the limit-of-detection of this defect mode, placed in the center of a PC that is 11-unit-cells in width as shown in Figure 2, is a modest $\delta t^{\text{lim}}/a=0.028$, which is the thinnest analyte coating thickness required to shift the resonance by one FWHM. For the unfragmented domain-wall defect of Figure 4, placed in the center of a photonic crystal that is 13-unit-cells in width, the sensitivity is improved to a frequency shift of $0.000665$ $2\pi c/a]$ per analyte-thickness $t=0.05a$ and the limit-of-detection is improved to $\delta t^{\text{lim}}/a=0.0023$. Since the quality factor increases exponentially with the number of unit cells of the PC, a corresponding exponential decrease in $\delta t^{\text{lim}}/a$ will occur.
@@
[…]
For our enhanced sensitivity chip (in Fig. 5) with fragmented rectangular strip construction, the limit-of-detection is greatly improved even for the narrower photonic crystal chip of 11-unit-cells in width. With such a modified domain-wall defect with resonant frequency $\omega=0.296732[2\pi a/c]$ and quality factor $Q=1312.8$, $\Delta\omega^{\text{FWHM}}=2.3\times 10^{-4}[2\pi a/c]$, and the limit-of-detection is $\delta t^{\text{lim}}/a=0.002$ in terms of analyte coating thickness. If the photonic crystal has a lattice constant of $a=1$ micron, a two-nanometer coating of analyte (of refractive index 1.45) would be detected. This minimal detectable analyte thickness can be further improved by increasing the number of unit cells of the photonic crystal as discussed above. For the 11-unit-cell-wide photonic crystal, the minimum detectable refractive index change of the background fluid is $\delta n^{\text{lim}}=7.35\times 10^{-4}$RIU. Again, this can be improved by increasing the number of photonic crystal unit cells in the light propagation direction.
@@ III Two topological defects: mode hybridization and Spectral independence
For biosensors capable of multi-parametric detection, which can distinguish different concentrations of multiple disease markers in a single measurement, multiple defects are required. This gives rise to mode hybridization, which complicates the interpretation of transmission spectra due to correlated spectral responses from different analyte-bindings. We first study a chip with two defects to analyze mode hybridization and its relation to transmission-levels. We analyze correlations in frequency shifts of the two separate resonances, and how to achieve spectral independence.
@@ III.3 Intermediate Hybridization
We functionalize this chip for biosensing and numerically simulate its transmission spectra for the various analyte-binding cases, as shown in Fig. 10(b)-(d). Due to increased mode hybridization, the spectral responses of the two resonances become more correlated. Nevertheless, different binding cases are clearly distinguishable especially if we monitor the transmission-level changes, as shown in the Truth Table (Fig. 10(e)). The frequency-shifts by themselves are also distinct. For example, with L-binding, the left-like resonance $|\text{L}\rangle^{\prime}$ will red-shift, inducing a correlated red-shift of the right-like resonance $|\text{R}\rangle^{\prime}$. Since the field strength of the left-like mode is higher than that of the right-like mode in the left-defect region, the left-like resonance $|\text{L}\rangle^{\prime}$ will red-shift more. The resonant frequency difference $\Delta\omega$ becomes larger with L-binding. This reduces hybridization and makes the transmission-levels of both resonances decline. Similar analysis applies to the other binding cases.
@@ V Conclusion and Outlook
Even higher sensitivity can be achieved with more elaborate designs or other types of defects. For example, increasing the number of rectangular strips within one unit cell from eight to twelve, with properly chosen strip width, such as $w_{def}=0.02a$, will improve the sensitivity further to 0.00758 $2\pi c/a]$ frequency shift per analyte-thickness $t=0.05a$, which is $46\%$ higher than our optimal sensitivity here. However, such a chip may be more challenging to fabricate. The chips given here all involve topological defects that involve a disruption of lattice periodicity. These were combined with substitutional defects (involving a shape change within a unit cell) to enhance sensitivity and adjust mode frequencies. Purely substitutional defect with fragmented rectangular strips can also achieve a high sensitivity. However, such chips typically require period doubling along the direction of fluid flow to enable optical coupling by an external plane wave.Abdullah15 ; Shuai16 ; Abdullah19 ; Dragan21 Combining topological domain-walls with substitutional defects enables additional design flexibility, not possible with purely substitutional defects. This may be useful in the design of 3D chips with finite pillar height. Nevertheless, other issues will arise. For example, an improperly fragmented unit cell may radiate more light into the 3rd dimension, thereby lowering the quality factor. Accordingly, further design considerations may be required to retain high sensitivity while not not losing light vertically.
There are specific types of fabrication errors that may be consequential to device functionality. One type is surface roughness or random shape fluctuations of the silicon structures (depicted in our manuscript as black squares and rectangles). The PBG of our structures provides considerable robustness against such fabrication errors. In our 2D simulation, these simply cause some small shift in the frequencies of the optical transmission resonances. This is a matter of calibration of each fabricated device. Although the transmission resonances may vary slightly from one device to another, the shift in resonance frequencies due to analyte binding is the same from one device to another. In other words, the user should simply calibrate the analyte-induced spectral fingerprint with respect to the spectral fingerprint in the absence of analyte binding.
In a real-world, 3D biosensor, the black squares and rectangles representing silicon must be vertically extruded to describe finite-height silicon pillars. A very consequential fabrication error is the possibility of shape fluctuations of the pillars in the vertical direction. This can lead to scattering of light (loss) into the vertical direction and consequent reduction of the in-plane transmission resonance quality factors. This is particularly important if there are very thin rectangular strips of silicon as seen in our fragmented defects (Fig. 5). This is an important topic for future research.
We hope that our design guidelines described herein will aid the development of real-world 3D photonic crystal biosensors, requiring minimal biofluid volume, while achieving low limit-of-detection, high sensitivity, and spectral independence, with efficacious transmission quality.